Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: C config; 1.06 OPT has some issues with subband frequency organization (limited observing opportunities over the weekend).
  • Key dates:
  • Staff Travel:
  • Visitors:
    • Brent Carlson: 18 October - 25 October; second trip post SKA meeting TBD

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
  • WIDAR (Rupen):
    • Post-Brent visit activities pushing on new features for commissioning tests (fast dumps - testing using 'funny' dump times (non-multiple of a second, e.g., 251 microseconds) - to solve issues with potential beating with the correlator dump times and the fringe), new features to be deployed (recirculation). Bob Sault visiting next week; will again tackle amplitude variations.
  • Computing (Robnett):

Software status and updates (Butler)

  • Links:
  • Notes: Some issues with OPT 1.06 release. New release of TelCal this week (minor improvements in robustness and flag handling).
  • CASA Update (new functionality available/issues/enhancements):Jeff noted a delay in the release due to a very slow split in ALMA data; this seems to be resolved and should be the last patch (split as cp should be much faster; averaging is currently unchanged).
  • AIPS Update (new functionality available/issues/enhancements): Rick noted some confusion in the reference channel after using the OBIT filler.
  • SDM Update:
    • Waiting on SDM/MCAF update to CalDevice table.

Observing recommendations update (Ops)

Commissioning priorities for the week (McMullin, Chandler)

Single maintenance day week:
  • Software updates (CBE/CM testing; focused on robustness; recirculation, faster dump times for testing)
  • Science
  • Other commissioning tests
  • Interrupt status for additional 3-bit testing

Commissioning Reports

Early Science 2011: McMullin/Rupen: Proposed guidelines for 2011 D configuration

  • OSRO
    • Up to 2 GHz bandwidth, full pol'n
    • Doppler setting (NOT tracking) independently for each baseband
    • spectral points as narrow as 0.122 kHz
    • In particular:
      • Two independent basebands, each with 8 subband pairs
        • The two basebands are independently tunable, but all 8 subbands within a given baseband must be adjacent
      • All subbands in both basebands must have the same bandwidth
      • Choice of full, dual, or single polarization products
        • 64 channels per sb per pol'n product for full pol'n
        • 128 channels per sb per pol'n product for dual pol'n
        • 256 channels per sb per pol'n product for single pol'n
      • No Doppler tracking
      • This gives 13.7 MB/dump, vs. current OSRO 1.7 MB/dump.
        • For D,C,B configurations: set min integration time of 3 sec (5 MB/s, 18 GB/hour, 422 GB/day, 50 TB/120 days)
        • For A configuration: set min integration time of 1 sec (13.7 MB/s, 48 GB/hour, 1.1 TB/day, 135 TB/120 days)
  • RSRO (in the absence of 3-bit samplers and 8 GHz BW):
    • Up to 2 GHz bandwidth, full pol'n; but with additional features for resident scientists:
      • Two independent basebands, each with more than 8 subband pairs; they may have different number of polarization products and different BWs.
      • Subbands are individually tunable (for access to widely spaced spectral lines)
      • Subbands may have different BWs.
    • Choice of full, dual, or single polarization products
      • 64 channels per sb per pol'n product for full pol'n
      • 128 channels per sb per pol'n product for dual pol'n
      • 256 channels per sb per pol'n product for single pol'n
    • Doppler tracking

General: Data Review

  • Antenna(s) - antennas (or baselines) on which the issue is seen.
  • Band(s) - seen just at one band or more
  • Polarization(s) - see in one or both polarizations
  • IF(s) - which IFs (A, B, C, D)
  • Date
  • Description of issue
  • Relevant JIRA - for reviewing the plots/additional info
  • Characteristic (behavior) - e.g., always, periodic, occasional, etc
  • SB name from archive
  • Any other comments (really helps to include baseband centers/LOs if possible).

Solar: Perley: Update from observations: Observations of sun with no attenuation in LSC converter (but adjustments in T304 to keep samplers in range)-get fringes! We can handle the 'quiet' sun. Work in executor to fix the GR correction in/out. Next test will at 20 dB attenuator in LSC.

Calibration: Perley: Switched Power update from testing: Vivek has been analyzing the 1 dB in/out; switched power changes by the expected ration; application restores power as hoped. Still checking level of accuracy (requirements are 1%!). Setting up tests using Cyg A (4x sys temps); hold attenuator fixed and let it vary.

High Frequency: Perley: Reference Pointing monitoring update: set of tests to compare CXK pointing at multiple bands; X band working fine; some antenna specific oddities.

Group reports

In preparation

  • Polarization: Multi-band leakage/PA (Myers, Marvil): Not a priority but still needed; keep on list.
  • Correlator: Krauss/Hunter: 7 vs 8 node observing (follow-up; pending RSRO results)
  • Array: Dhawan/Rupen: RF switch tests with rapid monitoring

Commissioning Test Planning

  • Spectral Line: Momjian: Zeeman suite of tests.
  • Polarization:
  • OPT: webtest of 1.06 capabilities/bug fixes (*); Deployed today.
  • Correlator: CBE averaging with restricted nodes
  • Calibration: Switched power tracking over several airmasses.
  • Polarization: Antenna cross polarization (Perley; Need script.); LSCX on 3C147; Scheduled for double maintenance day this week. Postponed until wideband X available on EA24
  • Spectral Line: Dhawan: Follow-up on frequency offset phase jumps characterization.
  • Spectral Line: Ott: Doppler tracking sampling follow-up
  • Polarization: Testing outline

RSRO/ECSO Reports

10B-183: Marvil update; updates in antenna/IF oddities (noted in JIRA ticket(s).

10B-209: Hallinan update; will redo test observation this week.

AC982: Chandler update; oddities with a suite of antennas; under review. Will present early results at AAS.

10B-154: Owen update; still on initial inspection.

Other updates:

  • Notes:
    • Current Residents
      • Satoko Takahashi (POC: Juergen Ott); area(s): continuum, mosaicing; 10B-126 (AT380): A Millimeter Study of the Embedded Star-Cluster in the Orion Molecular Cloud.
        • Hauyu Liu (POC: Juergen Ott); area(s): continuum, mosaicing
      • Greg Hallinan (POC: Kumar Golap); area(s):; 10B-209 (AH1020): Broadband Periodic Dynamic Spectra of Ultracool Dwarf Pulsars
      • Adam Leroy (POC: TBD); area(s):; 10A-196 (AL746): Resolving the Starbursts in Nearby LIRGs and ULIRGs
      • Jill Knapp (POC: Rupen/Claussen); area(s): recirculation; IRC+10216 demo science; 10B-200;10B-186. Welcome again.

Demo Science (Chandler/Condon): Project Codes: *TDEM*

Commissioning Group updates

  • OSRO User Support (van Moorsel): Link: https://staff.nrao.edu/wiki/bin/view/EVLA/OSROTests
    • Notes:
  • Observing Calibration/Flagging Strategies:
    • High Frequency (Claussen/Sjouwerman):
      • Notes:
    • Mid Frequency (Dhawan):
      • Notes:
    • Low Frequency (Momjian):
      • Notes:
  • Observing Modes
  • Calibrator Models
    • Notes:
  • HPC (Bhatnagar/Owen):
    • Notes:
  • Algorithm (Bhatnagar, Owen):
    • Notes: Discussion by Sanjay; note the Cyg A data reduced by LOFAR (found to be better); did imaging with multiscale clean and get the same quality though still not at the thermal noise limit (by a factor of a 1000); dynamic range is 5-8K. Wideband imaging (MS-MFS tests using AB1345; progress understanding spectral index errors for extended emission; will combine with primary beam corrections (good for narrow band; not so much for wide). Using autoflagging (demo in two weeks; show off at lunch tomorrow). Myers noted the first tests with OTF Mosaicing.

AOB

  • Frazer noted that the OPT/OST does not readily allow restrictions based on proximity to the sun. In addition, the current OST weather trending (phase for 10 minutes, wind for 2 hours) used for determining viability needs work. This will be kicked to the scheduling group for discussion/resolution. Some could be requested for development of 1.07. -> Action
  • Jeff noted that ALMA uses the the terminology of
    • Single spectral region modes (all subbands are contiguous);
    • Multi-spectral region modes (subbands can be independently tuned but must have the same number of channels)
    • Multi-spectral resolution modes (subbands can have different numbers of channels).

Misc Actions

  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin/Durand]: Clean up outdated MainSaver issues.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

-- JosephMcMullin - 2010-11-01
Topic revision: r3 - 2010-11-08, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback